Tele-Reco is a new specialty store that sells mobile phones, laptops and other computer accessories. A new Japanese-manufactured mobile phone reorder costs Tele-Reco $600 per unit. Tele-Reco’s annual holding cost rate is 22%. Ordering costs are estimated to be $70 per order.
1
Expert's answer
2020-06-15T19:22:44-0400

If the daily demand is expected to be 20 units monthly, the recommended order quantity will be:

"\text{Q}=\sqrt{\dfrac{2DO}{H}}"


The annual demand will be: "20\u00d712=240"

Ordering cost is $70.

The holding cost is: "\dfrac{22}{100}\times600=132"


"\text{Q}=\sqrt{\dfrac{2\times 240 \times 70}{132}}"


"\text{Q}=15.95=16 \text{Units}"